Spectral and Scattering Theory and Related Fields

in Research Institute of Mathematical Sciences of Kyoto University

December 4, Monday

13:30–14;10 Shu Nakamura

L^{p}-Theory of spectral shift function and the Wigner estimate 14:20–15:00 Kazushi Yoshitomi

Eigenvalue problems on domains with cracks 15:20–16:00 Atsushi Katsuda

BC-method and stability of Gel’fand inverse spectral problem 16:10–16:50 Kiyoshi Mochizuki

Jäger–Rejto approach on growth estimates

December 5, Tuesday

9:30–10:10 Mitsuteru Kadowaki

On existence of scattering solutions for dissipative systems 10:20–11:00 Fumio Hiroshima

Gibbs measures and a boson localization of a scalar quantum field model

11:10–11:50 Ingrid Beltita

On an abstract radiation condition 13:30–14:10 Arne Jensen

A unified approach to resolvent expansions at thresholds 14:20–15:00 Tadayoshi Adachi

Scattering theory for N-body quantum systems in a time-periodic electric field

15:20–16:00 Hiroshi Isozaki Asymptotic properties of solutions to 3–particle Schrödinger equations

16:10–16:50 Takashi Ichinose and Hideo Tamura On the norm convergence of the Trotter–Kato product formula with error bound

December 6, Wednesday

9:30–10:10 Kenji Nakanishi Nonrelativistic limit of scattering theory for nonlinear Klein-Gordon equations 10:20–11:00 Kenji Yajima and Gouping Zhang

Smoothing property of Schrödinger equations with potentials superquadratic at infinity

11:10–11:50 Hiroshi Ito and Hideo Tamura Scattering by magnetic fields at large separation in two dimensions

13:20–14:00 Wataru Ichinose The Feymann path integral representation of Green functions of position and momentum operators

14:10–14:50 Takuya Mine On negative eigenvalues of Schrödinger operators in magnetic fields with Robin boundary conditions

15:00–15:40 Takashi Ōkaji Absence of eigenvalues of Maxwell operators